OpenStudy (anonymous):

Please, solve the question below. It's a tad complicated so I'm going to have to use the equation generator here to show it to you. When you get your answer, could you please show me the steps. Thank you!

OpenStudy (anonymous):

Here's the question. Add \[\frac{ 2 }{ e-3 }+\frac{ 2 }{ e^2-9 }\] Simplify if possible.

OpenStudy (anonymous):

And here are my possible choices. A.\[\frac{ 2(e-2) }{ (e-3)(e+3) }\] B. \[\frac{ 4e }{ (e-3)(e+3) }\] C. \[\frac{ 2(e+4) }{ (e-3)(e+3) }\] D. \[\frac{ 4 }{ (e-3)(e+3) }\]

OpenStudy (phi):

you need a common denominator. It helps if you can factor (e^2-9) (a difference of squares) can you factor e^2 - 9 ?

OpenStudy (anonymous):

@phi I think I got it! Is it C? I'm just worried that that "e" is representing the base of exponential function. But would it still be just like simplifying rationals?

OpenStudy (anonymous):

C is the correct answer

OpenStudy (phi):

It does not really matter what e stands for. I did not do the problem, but the first thing to do is find the common denominator. one way to find the common denominator is factor e^2-9 to see what factors is has.

OpenStudy (anonymous):

Thank you so much! You have been wonderful!